# AstroMLab 3: Achieving GPT-4o Level Performance in Astronomy with a Specialized 8B-Parameter Large Language Model

### v1

@@ Abstract
AstroSage-Llama-3.1-8B is a domain-specialized natural-language AI assistant tailored for research in astronomy, astrophysics, and cosmology. Trained on the complete collection of astronomy-related arXiv papers from 2007-2024 along with millions of synthetically-generated question-answer pairs and other astronomical literature, AstroSage-Llama-3.1-8B demonstrates remarkable proficiency on a wide range of questions. AstroSage-Llama-3.1-8B scores $80.9\%$ on the AstroMLab-1 benchmark, greatly outperforming all models—proprietary and open-weight—in the 8-billion parameter class, and performing on par with GPT-4o. This achievement demonstrates the potential of domain specialization in AI, suggesting that focused training can yield capabilities exceeding those of much larger, general-purpose models. AstroSage-Llama-3.1-8B is freely available, enabling widespread access to advanced AI capabilities for astronomical education and research.
@@ I Introduction
Building on the previous efforts of cosmosage [8] and AstroLLaMA, we have developed AstroSage-Llama-3.1-8B, a natural language assistant specialized in astronomy, astrophysics, cosmology, and astronomical instrumentation. For the remainder of this paper, we will refer to these subdomains collectively as “astronomy”. Through the use of a substantially more extensive and well-curated training dataset, we demonstrate for the first time that our specialized language model significantly outperforms baseline models in downstream tasks, particularly in astronomical knowledge recall.
In the long term, we envision an agentic research assistant capable of autonomously conducting literature reviews, identifying relevant hypotheses, carrying out data analysis, and even formulating new research questions. The development of such scientific agents (LLMs capable of solving scientific problems end-to-end) is already a rapidly growing field in astronomy. Recent studies have shown promising results in automating research tasks, such as analyzing James Webb Space Telescope data through multi-agent collaboration and self-play reinforcement learning [9]. However, these studies have been largely constrained by the substantial API costs associated with proprietary models.
@@ II.1 Dataset Preparation
  • Approximately 250,000 arXiv preprints from 2007-2024 with primary or cross-listed categories in astro-ph (astrophysics) or gr-qc (general relativity and quantum cosmology). We deliberately excluded the Annual Review of Astronomy and Astrophysics papers used in Ting et al. [11, AstroMLab-1] to generate the benchmark questions, ensuring our evaluation would test the model’s ability to generalize knowledge rather than recall specific source materials.
  • Relevant articles from a depth-2 search through Wikipedia’s astronomy and astrophysics categories.
  • A selection of textbooks that are available as PDFs or ebooks online.
We processed the data into markdown format. For the vast majority of the dataset, the rendered PDF files were converted to markdown using Nougat OCR [16]. For the remaining sources, the data was either already in markdown format, or was left as plain text.
@@ II.2 Pretraining run
The Llama-3.1 tokenizer, which uses a variant of tiktoken with UTF-flexible encoding, was sufficient for our purposes, so we did not introduce any astronomy-specific tokens to the vocabulary. We considered incorporating arXiv identifiers (in the format arXiv:YYMM.numbervV) for papers in the training set but ultimately decided against it due to the substantial increase in vocabulary size this would entail. Future work may explore expanding the vocabulary to include numerical representations relevant to astronomy, such as common quantities and units.
During training, we tracked the loss function and step sizes. Minimal hyperparameter tuning was performed, relying on the hyperparameters from de Haan [8] for all parameters other than the learning rate. The learning rate was extrapolated from smaller runs, but remained problematic, as Frontier requires high levels of parallelization with short wall times. This caused our initial runs to suffer from either insufficient learning due to a low learning rate, or catastrophic exploding gradients due to an excessively high learning rate. The final run described in this work used a tuned learning schedule with a learning rate as high as possible but still allowing convergence. In future efforts, we plan to further optimize the efficiency of the training procedure along the lines of the work presented in Dash et al. [17], incorporating tensor, pipeline, and data paralellism through libraries such as Megatron-Deepspeed. We also aim to request a longer walltime with lower parallelization factor in order to be able to reduce loss and improve downstream performance.
@@ II.3 Dataset Cleaning
Based on the distribution of perplexity scores (Figure 1), we established a threshold that excluded the top 2% of paragraphs with the highest perplexity scores. We then reconstructed each document using only the paragraphs below this threshold. This cleaning procedure removed approximately 2% of the total data volume. Figure 1 illustrates this process by showing the distribution of perplexity scores and our chosen threshold.
@@ III.1 SFT Dataset
Pan et al. [7] identified a critical limitation in the AstroLLaMA series of specialized astronomical LLMs: their inability to outperform even their own starting base models, partly due to inadequate SFT. While these models showed marginal improvements in basic next-token prediction tasks, they performed worse than their baseline models on instructional Q&\&&A tasks, even for straightforward astronomical knowledge recall. This shortcoming fundamentally undermines the purpose of specialized training. Therefore, in our study, we paid particular attention to the SFT process, generating training datasets orders of magnitude larger than previously available in astronomy.
The largest component of our SFT dataset consists of question-and-answer (Q&\&&A) pairs. Using the method from de Haan [8], we generated over 11 million synthetic Q&\&&A pairs from papers in our CPT dataset. These Q&\&&A pairs were then evaluated using an LLM based on four criteria:
@@
    Stand-alone: The ability of each Q&\&&A pair to be understood in isolation, without needing additional context beyond the content provided. This ensures that an expert in astronomy, astrophysics, or cosmology could answer the question based solely on the information within the pair.
@@
    Pertinence: The relevance and importance of the question to researchers or students in astronomy, astrophysics, or cosmology. We want the questions to be ones that a professional in the field might ask or find valuable for deeper understanding.
@@
    Overall Quality
The LLM was presented with several hand-written examples of ratings in the form Question–Answer–Score–Explanation, followed by the Question–Answer pair to be judged. The model was then asked to complete the score values. A small number of resulting scores were verified and confirmed to be sufficiently accurate. Only Q&\&&A pairs with perfect scores in all four categories were kept, resulting in 8.8 million high-quality Q&\&&A pairs.
@@
Furthermore, we generated a metadata-based dataset, where a again a series of custom rules and random selections result in diverse questions about titles, dates of publication, arXiv IDs, and first author names from the papers in the CPT dataset. This was included in an effort to memorize the paper metadata so that users can reference papers in their conversations with AstroSage-Llama-3.1-8B. However, as we will discuss, our training procedure was relatively shallow and insufficient memorization of this information took place, as spot checks show that the final model can not reliably answer metadata-based questions correctly.
@@ III.2 SFT procedure
Figure 2: Supervised fine-tuning loss curve. The learning rate schedule, shown in brick red, follows a quadratic warmup followed by a cosine schedule that ends at 10% of the peak learning rate. The peak learning rate was chosen to prevent exploding gradients. The loss curve in black shows no significant signs of overfitting, as evidenced by minimal discontinuities at epoch boundaries. The green curve represents the L2 norm of the parameter update, shown in arbitrary units.
@@ IV Model Merging
Figure 3: Performance comparison on the AstroMLab 1 benchmark, which contains 4,425 high-quality, human-verified multiple-choice questions across astronomy, astrophysics, cosmology, and instrumentation. The gray shaded region indicates the performance range of human domain experts. We present updated results as of November 2024, incorporating both cutting-edge proprietary and open-weight models. AstroSage-Llama-3.1-8B outperforms all other models in the 8-billion parameter class and achieves performance comparable to OpenAI’s latest models, including GPT-4o, while Claude-3.5-Sonnet maintains the highest performance overall. The diagonal dashed lines represent cost-efficiency trade-offs as determined in AstroMLab 1 (see text for details). The Wilson Score interval shows the typical uncertainty in the score due to the finite number of questions. Star symbols indicate all published specialized LLMs for astronomy to our knowledge. Previously, these specialized models often failed to outperform their baseline models in astronomical recall due to various training limitations. AstroSage-Llama-3.1-8B represents a significant advancement in specialized astronomical LLMs, demonstrating that extensive data curation, massive continued pre-training and supervised fine-tuning, and model merging techniques can substantially improve performance on specific astronomical tasks. This result highlights the effectiveness of domain specialization even in relatively smaller models.
@@ V Evaluation
Figure 4: Performance comparison across general language model benchmarks. The left panel shows a bar chart comparing model performances on standard benchmarks. The center radar chart visualizes the effect of continued pretraining. The right radar chart compares our model post-SFT against BAAI/Infinity-Instruct-7M-Gen-LLaMA3_1-8B, which was trained on the same base model and SFT data but without our astronomy-specific data. Despite optimization for astronomy tasks, the merged model we are releasing as AstroSage-Llama-3.1-8B maintains strong general capabilities in reasoning, mathematics, and coding, demonstrating that domain specialization did not come at the cost of other abilities.
@@
The performance score is calculated as the fraction of correctly answered multiple-choice questions in the benchmarking dataset. The resulting scores are shown in Figure 3, where round symbols represent scores for cutting-edge proprietary and open-weight models. The open-weight models are also marked with an outer circle. The x-axis displays the cost per $10^{5}$ tokens, a metric chosen based on practical applications: in the first (and to our knowledge, only) implementation of astronomical agents [9], analyzing a celestial source’s spectral energy distribution from James Webb Space Telescope data requires approximately $10^{5}$ tokens. The top x-axis shows costs scaled to 3B ($3\times 10^{9}$) tokens, roughly equivalent to the entire astro-ph section of the arXiv. For proprietary models, we use current token costs (averaging input and output costs where they differ), while open-weight model costs are estimated based on typical pricing of commercial API platforms.
@@
To establish a human performance baseline, two domain experts from our team independently completed a random subset of benchmark questions under controlled conditions. Each expert was allowed approximately 30 seconds per question and prohibited from consulting external references, including web searches or language model assistance. Both experts achieved remarkably consistent scores of approximately 68%, which we designate as the “human domain expert baseline.” The fact that most evaluated LLMs significantly surpassed this baseline demonstrates both the benchmark’s comprehensive scope and difficulty, while highlighting the remarkable capabilities of current LLMs in capturing and applying complex astronomical knowledge.
@@
To ensure our domain specialization didn’t compromise general capabilities, we evaluated AstroSage-Llama-3.1-8B across a comprehensive suite of standard language model benchmarks. These include IF-EVAL (instruction following), BBH (binary hypothesis testing), MATH (mathematical reasoning), GPQA (graduate-level science questions), MUSR (real-world decision-making scenarios), and MMLU-PRO (an expanded version of MMLU with more challenging reasoning questions). As shown in Figure 4, our CPT+SFT model (green, initialized from the Llama-3.1 base model) initially performed below the Llama-3.1 instruct model (purple) on five out of the six non-astronomy benchmarks. This was expected, given that Meta’s proprietary SFT dataset for their instruct model likely far exceeds what’s feasible for an academic research group to reproduce. The merging procedure, pulling in only 25% of its weight from Meta-Llama-3.1-8B-Instruct, allowed us to recover much of this performance deficit.
Crucially, this performance recovery through model merging did not compromise AstroSage-Llama-3.1-8B’s astronomical expertise—it maintained its 8-point improvement (representing more than 100-fold increase in cost-effectiveness) on astronomical Q&\&&A tasks while largely preserving capabilities across most general benchmarks. The only notable performance decrease occurred in IF-EVAL, which tests instruction following. This limited decline is unsurprising, as instruction following remains one of the more brittle capabilities in language models and likely heavily depends on the proprietary training data used in Meta’s instruct model. In fact, when compared AstroSage-Llama-3.1-8B to BAAI/Infinity-Instruct-7M-Gen-LLaMA3 1-8B, the latter shows an even more severe performance deficit, highlighting how our refined training strategy and expanded SFT dataset represent crucial improvements. Ultimately, our model merging approach successfully preserved most general capabilities without sacrificing the gained astronomical expertise. This balance is essential, as it enables AstroSage-Llama-3.1-8B to engage in natural conversations and assist with broader tasks while excelling in astronomy-specific applications.

### v2

# Achieving GPT-4o Level Performance in Astronomy with a Specialized 8B-Parameter Large Language Model

@@ Abstract
AstroSage-Llama-3.1-8B is a domain-specialized natural-language AI assistant tailored for research in astronomy, astrophysics, cosmology, and astronomical instrumentation. Trained on the complete collection of astronomy-related arXiv papers from 2007-2024 along with millions of synthetically-generated question-answer pairs and other astronomical literature, AstroSage-Llama-3.1-8B demonstrates remarkable proficiency on a wide range of questions. AstroSage-Llama-3.1-8B scores $80.9\%$ on the AstroMLab-1 benchmark, greatly outperforming all models—proprietary and open-weight—in the 8-billion parameter class, and performing on par with GPT-4o. This achievement demonstrates the potential of domain specialization in AI, suggesting that focused training can yield capabilities exceeding those of much larger, general-purpose models. AstroSage-Llama-3.1-8B is freely available, enabling widespread access to advanced AI capabilities for astronomical education and research.
@@ I Introduction
Building on the previous efforts of cosmosage [8] and AstroLLaMA, we have developed AstroSage-Llama-3.1-8B, a natural language assistant specialized in astronomy, astrophysics, cosmology, and astronomical instrumentation. For the remainder of this paper, we will refer to these subdomains collectively as “astronomy”. Through the use of a substantially more extensive and well-curated training dataset, we demonstrate for the first time that our specialized language model significantly outperforms baseline models in downstream tasks.
In the long term, we envision an agentic research assistant capable of autonomously conducting literature reviews, identifying relevant hypotheses, carrying out data analysis, and even formulating new research questions. The development of such scientific agents (LLMs capable of solving scientific problems end-to-end) is already a rapidly growing field in astronomy. Recent studies have shown promising results in automating research tasks, such as analyzing James Webb Space Telescope data through multi-agent collaboration and self-play reinforcement learning [9]. However, these studies have been largely constrained by the substantial costs associated with utilizing proprietary models at high volume.
@@ II.1 Dataset Preparation
  • Approximately 250,000 arXiv preprints from 2007-2024 with primary or cross-listed categories in astro-ph (astrophysics) or gr-qc (general relativity and quantum cosmology). For preprints with multiple versions, we selected the most recent. We deliberately excluded the Annual Review of Astronomy and Astrophysics papers used in Ting et al. [11, AstroMLab-1] to generate the benchmark questions, ensuring our evaluation would test the model’s ability to generalize knowledge rather than recall specific source materials.
  • Relevant articles from a depth-2 search through Wikipedia’s astronomy and astrophysics categories. We include articles that directly belong to these categories or their immediate subcategories, then further add linked articles. We perform keyword-based exclusion on a few tens of custom keywords such as “deity”, “cyberspace”, “mythological”, “sports” and so on, yielding a total of nearly 30,000 articles.
  • A selection of $\sim$800 textbooks that are available as PDFs or ebooks online.
For the vast majority of the dataset, the rendered PDF files were converted to Mathpix Markdown using Nougat OCR [16]. For the remaining sources, the data was either already in Markdown format, or was left as plain text.
@@ II.2 Pretraining run
The Llama-3.1 tokenizer, which uses a variant of tiktoken [17] with UTF-flexible encoding, was sufficient for our purposes, so we did not introduce any astronomy-specific tokens to the vocabulary. We considered incorporating arXiv identifiers (in the format arXiv:YYMM.numbervV) for papers in the training set but ultimately decided against it due to the substantial increase in vocabulary size this would entail. Future work may explore expanding the vocabulary to include numerical representations relevant to astronomy, such as common quantities and units.
During training, we tracked the loss function and step sizes. Minimal hyperparameter tuning was performed, relying on the hyperparameters from de Haan [8] for all parameters other than the learning rate schedule. The warmup period of 40 samples was determined empirically from the loss curve of an earlier, canceled CPT run. The peak learning rate was extrapolated from smaller runs, but remained problematic, as Frontier requires high levels of parallelization with short wall times. This caused our initial runs to suffer from either insufficient learning due to a low peak learning rate, or catastrophic exploding gradients due to an excessively high peak learning rate. The final run described in this work used a tuned learning schedule with a peak learning rate as high as possible but still allowing convergence. In future efforts, we plan to further optimize the efficiency of the training procedure along the lines of the work presented in Dash et al. [18], incorporating tensor, pipeline, and data paralellism through libraries such as Megatron-Deepspeed [19, 20]. We also aim to request a longer walltime with lower parallelization factor in order to be able to reduce loss and improve downstream performance.
@@ II.3 Dataset Cleaning
Based on the distribution of perplexity scores (Figure 1), as well as manual inspection of some examples, we established a threshold that excluded the top 2% of paragraphs with the highest perplexity scores. We then reconstructed each document using only the paragraphs below this threshold. This cleaning procedure removed approximately 2% of the total data volume. Figure 1 illustrates this process by showing the distribution of perplexity scores and our chosen threshold.
@@ III.1 SFT Dataset
Pan et al. [7] identified a critical limitation in the AstroLLaMA series of specialized astronomical LLMs: their inability to outperform even their own starting base models, partly due to inadequate SFT. While these models showed marginal improvements in basic next-token prediction tasks, they performed worse than their baseline models on instructional question-and-answer (Q&\&&A) tasks, even for straightforward astronomical knowledge recall. This shortcoming fundamentally undermines the purpose of specialized training. Therefore, in our study, we paid particular attention to the SFT process, generating training datasets orders of magnitude larger than previously available in astronomy.
The largest component of our SFT dataset consists of Q&\&&A pairs. Using the method from de Haan [8], we generated over 11 million synthetic Q&\&&A pairs from papers in our CPT dataset. These Q&\&&A pairs were then evaluated using an LLM based on four criteria:
@@
    Stand-alone: The ability of each Q&\&&A pair to be understood in isolation, without needing additional context beyond the content provided. This ensures that an astronomy expert could answer the question based solely on the information within the pair.
@@
    Pertinence: The relevance and importance of the question to researchers or students in astronomy. We want the questions to be ones that a professional in the field might ask or find valuable for deeper understanding.
@@
    Overall Quality: a holistic assessment of the Q&\&&A pair. This criterion is intended to capture aspects that may not be explicitly covered by correctness, stand-alone completeness, or pertinence but contribute to the overall usefulness of the Q&\&&A pair.
The LLM was presented with several hand-written examples of ratings in the form Question–Answer–Score–Explanation, followed by the Question–Answer pair to be judged. The model was then asked to complete the score values. A small number of resulting scores were verified and confirmed to be sufficiently accurate. Only Q&\&&A pairs with perfect scores in all four categories were kept, resulting in 8.8 million high-quality Q&\&&A pairs. Below are a few examples.
  • Q: What are the three distinct phases of galaxy formation identified in the EAGLE simulation, and how are they primarily determined?
    A: The three distinct phases of galaxy formation identified are:
    1. 1.
      Disorganized growth in low-mass galaxies (stellar mass less than approximately $10^{9.5}M_{\odot}$), characterized by random stellar motions and dominated by in-situ star formation and mergers;
    2. 2.
      Evolution towards a disc-dominated morphology in intermediate-mass galaxies (stellar mass between $10^{9.5}M_{\odot}$ and $10^{10.5}M_{\odot}$), with bulge formation associated with mergers;
    3. 3.
      High-mass galaxies (stellar mass greater than approximately $10^{10.5}M_{\odot}$) experience a slowdown in in-situ star formation, transitioning to spheroidal morphologies driven by the accretion of ex-situ stars. These phases are primarily determined by stellar mass rather than redshift.
  • Q: What are the expected contributions of AGN feedback to the heating and dynamics of the hot ISM?
    A: AGN feedback is anticipated to play a significant role in regulating the cooling of the hot ISM, with energy released during accretion often exceeding starburst-driven heating and driving turbulence and outflows that influence the dynamics of the ISM.
  • Q: What are the applications of the Augmented Lagrangian Perturbation Theory (ALPT) in cosmology?
    A: ALPT can be used for setting initial conditions for N-body simulations, generating mock galaxy catalogs, analyzing cosmic structures, and reconstructing primordial density fluctuations for baryon acoustic oscillation studies.
  • Q: What are the three effects that anisotropies in the optical depth produce in the CMB?
    A: Anisotropies in the optical depth produce three effects in the CMB: (i) screening of the temperature and polarization fluctuations observed today by an overall factor of $\exp(-\tau(\hat{n}))$; (ii) Thomson scattering, producing new E-mode polarization from the local temperature quadrupole, with subdominant contributions to B-modes from reionization effects; and (iii) new temperature anisotropy generated from the radial motion of ionized bubbles relative to the observer.
@@
Furthermore, we generated a metadata-based dataset, where again a series of custom rules and random selections result in diverse questions about titles, dates of publication, arXiv IDs, and first author names from the papers in the CPT dataset. This was included in an effort to memorize the paper metadata so that users can reference papers in their conversations with AstroSage-Llama-3.1-8B. However, we find that after training, the model is unable to reliably answer metadata-based questions such as producing a first author name from an arXiv ID. This failure to memorize could be ameliorated in future training runs by increasing the learning rate or number of epochs.
@@ III.2 SFT procedure
Figure 2: Supervised fine-tuning loss curve. The learning rate schedule, shown as the dashed brick red curve, follows a quadratic warmup followed by a cosine schedule that ends at 10% of the peak learning rate. The peak learning rate was chosen to prevent exploding gradients. The loss curve in black shows no significant signs of overfitting, as evidenced by minimal discontinuities at epoch boundaries. The green curve represents the L2 norm of the parameter update, shown in arbitrary units.
@@ IV Model Merging
Figure 3: Performance comparison on the AstroMLab 1 benchmark, which contains 4,425 high-quality, human-verified multiple-choice questions across astronomy, astrophysics, cosmology, and instrumentation. The gray shaded bar spans the performance of the two human domain experts. We present updated results as of November 2024, incorporating both cutting-edge proprietary and open-weight models. Open-weight models are circled. AstroSage-Llama-3.1-8B outperforms all other models in the 8-billion parameter class and achieves performance comparable to OpenAI’s latest models, including GPT-4o, while Claude-3.5-Sonnet maintains the highest performance overall. The diagonal dashed lines represent cost-efficiency trade-offs as determined in AstroMLab 1 (see text for details). The Wilson Score interval shows the typical uncertainty in the score due to the finite number of questions. Star symbols indicate all published specialized LLMs for astronomy to our knowledge. Previously, these specialized models often failed to outperform their baseline models in astronomical recall due to various training limitations. AstroSage-Llama-3.1-8B represents a significant advancement in specialized astronomical LLMs, demonstrating that extensive data curation, massive continued pre-training and supervised fine-tuning, and model merging techniques can substantially improve performance on specific astronomical tasks. This result highlights the effectiveness of domain specialization even in relatively smaller models.
@@ V Evaluation
We have performed three types of quantitative evaluation. The first is an evaluation of AstroSage-Llama-3.1-8B on a multiple-choice question benchmark, the second on several general-purpose benchmarks, and the third direct preference testing against a similar model that was not specialized in astronomy.
V.1 Multiple-Choice Question Benchmark
Figure 4: Performance comparison across general language model benchmarks. The bar chart shows a comparison of model performances on standard benchmarks. The left radar chart visualizes the effect of continued pretraining. The right radar chart compares our model post-SFT against BAAI/Infinity-Instruct-7M-Gen-LLaMA3_1-8B, which was trained on the same base model and SFT data but without our astronomy-specific data. Despite optimization for astronomy tasks, the merged model we are releasing as AstroSage-Llama-3.1-8B maintains strong general capabilities in reasoning, mathematics, and coding, demonstrating that domain specialization did not come at the cost of other abilities.
@@ V.1 Multiple-Choice Question Benchmark
The performance score is calculated as the fraction of correctly answered multiple-choice questions in the benchmarking dataset. The resulting scores are shown in Figure 3, where round symbols represent scores for cutting-edge proprietary and open-weight models. The open-weight models are also marked with an outer circle. The x-axis displays the cost per $10^{5}$ tokens, a metric chosen based on practical applications: in the first (and to our knowledge, only) implementation of astronomical agents [9], analyzing a celestial source’s spectral energy distribution from James Webb Space Telescope data requires approximately $10^{5}$ tokens. The top x-axis shows costs scaled to 3B ($3\times 10^{9}$) tokens, roughly equivalent to the entire astro-ph section of the arXiv. For proprietary models, we use current token costs (averaging input and output costs where they differ), while open-weight model costs are estimated based on typical pricing of commercial GPU platforms.
@@
To establish a human performance baseline, two domain experts from our team independently completed a random subset of benchmark questions under controlled conditions. The experts answered on the order of one hundred questions each, taking around 30 seconds per question. No external references, web searches, or language model assistance were used. Both experts achieved remarkably consistent scores of approximately 68%, which we designate as the “Indicative Human Domain Expert” score. The fact that most evaluated LLMs significantly surpassed this baseline demonstrates both the benchmark’s comprehensive scope and difficulty, while highlighting the remarkable capabilities of current LLMs in capturing and applying complex astronomical knowledge.
@@
V.2 General-Purpose Benchmarks
To ensure our domain specialization did not compromise general capabilities, we evaluated AstroSage-Llama-3.1-8B across a comprehensive suite of standard language model benchmarks. These include IF-EVAL [25] (instruction following), BBH [26] (binary hypothesis testing), MATH [27] (mathematical reasoning), GPQA [28](graduate-level science questions), MUSR [29] (real-world decision-making scenarios), and MMLU-PRO [30] (an expanded version of MMLU with more challenging reasoning questions). As shown in Figure 4, our CPT+SFT model (green, initialized from the Llama-3.1 base model) initially performed below the Llama-3.1 instruct model (purple) on five out of the six non-astronomy benchmarks. This was expected, given that Meta’s proprietary SFT dataset for their instruct model likely far exceeds what is feasible for an academic research group to reproduce. The merging procedure, pulling in only 25% of its weight from Meta-Llama-3.1-8B-Instruct, allowed us to recover much of this performance deficit.
Crucially, this performance recovery through model merging did not compromise AstroSage-Llama-3.1-8B’s astronomical expertise—it maintained its 8-point improvement (representing more than 100-fold increase in cost-effectiveness) on astronomical Q&\&&A tasks while largely preserving capabilities across most general benchmarks. The only notable performance decrease occurred in IF-EVAL, which tests instruction following. This limited decline is unsurprising, as instruction following remains one of the more brittle capabilities in language models and likely heavily depends on the proprietary training data used in Meta’s instruct model. In fact, when comparing AstroSage-Llama-3.1-8B to BAAI/Infinity-Instruct-7M-Gen-LLaMA3 1-8B, the latter shows an even more severe performance deficit, highlighting how our refined training strategy and expanded SFT dataset represent crucial improvements. Ultimately, our model merging approach successfully preserved most general capabilities without sacrificing the gained astronomical expertise. This balance is essential, as it enables AstroSage-Llama-3.1-8B to engage in natural conversations and assist with broader tasks while excelling in astronomy-specific applications.
V.3 Human Blind Rankings
Simultaneously with de Haan [8], we performed a blinded preference ranking of AstroSage-Llama-3.1-8B against Meta-Llama-3.1-8B-Instruct. Fifteen questions were written about diverse areas of cosmology, spanning layperson to professional-level difficulty. Independent evaluators compared responses from AstroSage-Llama-3.1-8B and Meta-Llama-3.1-8B-Instruct, with both models receiving identical system prompts tailored to the cosmology context.
Responses were presented in randomized order to the evaluators, who rated their quality without knowledge of the source. The evaluation involved three evaluators, who preferred the AstroSage-Llama-3.1-8B answers in 73% of cases, reflecting a statistically significant preference for AstroSage-Llama-3.1-8B.